# Revisiting Graph-based Fraud Detection in Sight of Heterophily and Spectrum

### v1

@@ Abstract
Graph-based fraud detection (GFD) can be regarded as a challenging semi-supervised node binary classification task. In recent years, Graph Neural Networks(GNN) have been widely applied to GFD, characterizing the anomalous possibility of a node by aggregating neighbor information. However, fraud graphs are inherently heterophilic, thus most of GNNs perform poorly due to their assumption of homophily. In addition, due to the existence of heterophily and class imbalance problem, the existing models do not fully utilize the precious node label information. To address the above issues, this paper proposes a semi-supervised GNN-based fraud detector SEC-GFD. This detector includes a hybrid filtering module and a local environmental constraint module, the two modules are utilized to solve heterophily and label utilization problem respectively. The first module starts from the perspective of the spectral domain, and solves the heterophily problem to a certain extent. Specifically, it divides the spectrum into multiple mixed frequency bands according to the correlation between spectrum energy distribution and heterophily. Then in order to make full use of the node label information, a local environmental constraint module is adaptively designed. The comprehensive experimental results on four real-world fraud detection datasets show that SEC-GFD outperforms other competitive graph-based fraud detectors.
@@ Introduction
With the rapid development of the Internet, various fraudulent entiies emerge in fraud activities, and the huge losses caused by this have attracted continuous attention from industry, academia and state agencies(Wang et al. 2018; Pourhabibi et al. 2020; Liu et al. 2021a; Zhang et al. 2021). For instance, fraudulent behaviors aimed at online payment have caused huge property losses to clients; fraudulent or malicious comments on websites may bring negative effects on merchants and consumers.
@@
In recent years, graph-based fraud detection methods(Hooi et al. 2017; Qin et al. 2022) have been widely used in practical applications. Using entities in fraud scenarios as nodes and interactions between entities as edges, thus graph-level fraud detection can be achieved. Furthermore, based on the constructed fraud graph, as a deep neural network for graph representation learning, Graph Neural Network(GNN) has been demonstrated to be effective in this task. Specifically, GNN employs neural modules to aggregate neighbor information and update node representations, enabling the discrimination of fraudulent entities.
@@
However, fraud detection scenarios are inherently heterophily(Zhu et al. 2021), i.e., target nodes (like node 5 in Fig. 1) and their direct neighbors are prone to different class labels or features. In practical applications, fraudulent entities may use benign entities as springboards to perform high-order interactions with multi-hop neighbors to accomplish fraudulent actions or transmit fraudulent information(for example, fraud nodes 3 and 6 are two-hop neighbors in Fig. 1). For instance, in financial transaction networks(Akoglu, Tong, and Koutra 2015), fraudsters often use non-fraudulent users to conduct transactions.
@@
Due to the existence of heterophily, the performance of conventional graph neural network is poor. This is because vanilla GNN is naturally a low-pass filter, which forces the representation of adjacent nodes to be similar during the information aggregation process. Then it may destroy the discriminative information of anomaly nodes. Furthermore, within the domain of fraud detection, heterophily graph datasets exhibit highly imbalanced class distribution properties. This phenomenon results in anomaly nodes being submerged in normal nodes, which manifests as high homophily of normal nodes and high heterophily of anomaly nodes. This is also called the common ”camouflage” problem(Zhang et al. 2023) in fraud detection scenarios. Moreover, current semi-supervised fraud detection methods based on Graph Neural Networks (GNNs) suffer from limitations in terms of label utilization. These approaches typically learn node representations through a message passing mechanism, extract node embeddings from the final layer of the network, and incorporate label information during the training process. Evidently, these approaches overlook the intricate high-order associations between node label information, node characteristics, and their contextual surroundings.
@@
The current paramount concern entails addressing the heterophily issue within the context of extreme category imbalance. The advanced methods can be divided into two types. The first type of the methods is to design a new message aggregation mechanism to simultaneously or separately consider the characteristics of homophily and heterophily edges. And the second type of the methods is to identify the heterophily edges on fraud graphs and then perform pruning. According to recent researches, researchers found the spectral ”right shift” phenomenon(Tang et al. 2022) on both synthetic and real datasets, proving that the higher the anomaly ratio, the more high-frequency components on the graph spectrum. In addition, researchers also demonstrated that, when using the ground truth labels as signals, the proportion of heterophily edges is positively correlated with the spectral energy distribution(Gao et al. 2023a). Based on these, we conclude that heterophily is not universally harmful, instead it should respond distinctively when meeting different filters. In order to verify our conjecture, we conducted sufficient experiments, and the specific related discussions are placed in Experiments.
Figure 1: Schematic diagram of fraud relationship in real scenarios
@@
In order to design a GNN-based fraud detector that adapts to high heterophily and has remarkable label utilization, we first propose a conjecture that heterophily edge clipping responds differently to different filter functions, and conduct comprehensive experiments to verify the conjecture. Based on the results, we design a Spectrum-Enhanced and Environment-Constrainted Graph Fraud Detector (SEC-GFD) to comprehensively integrate rich spectrum and label information into a fraud detector. The framework of the proposed method is shown in Fig. 2. The SEC-GFD framework consists of two different modules that address the aforementioned two challenges respectively. The hybrid-pass filter module employs a sophisticated message passing technique utilizing the decomposed spectrum to acquire comprehensive high-frequency representations. The local environmental constraint module enhances the higher-order connections between nodes’ true labels and their surrounding environments, thereby mitigating the issue of limited label utilization. The main contributions of our method are summarized as follows:
@@
  • To address the challenge of heterophily, this paper employs spectral analysis to partition the graph spectrum into hybrid frequency bands. Then a novel hybrid spectral filter is designed to conduct message passing on these frequency bands separately and finally aggregated.
@@
Semi-supervised node classification is one of the most important tasks in graph learning. In recent years, GNNs have achieved excellent performance in semi-supervised node classification, like GCN(Kipf and Welling 2016), GraphSAGE(Hamilton, Ying, and Leskovec 2017), GAT(Veličković et al. 2017), GIN(Xu et al. 2018). Compared with traditional graph embedding methods such as DeepWalk(Perozzi, Al-Rfou, and Skiena 2014) and Node2vec(Grover and Leskovec 2016), which mainly focus on graph structure, GNN can consider both node features and graph structure. Each layer of GNN first performs message passing according to the node neighbors, and then calculates the node representation through the aggregated information.
Figure 2: The above image presents an overview of our model SEC-GFD, where Fig (a) and (b) respectively demonstrate the details of the hybrid-pass filter module and local environmental constraint module.
@@
GNN-based fraud detection(GFD) can be defined as an imbalanced binary classification task that identifies outlier nodes that deviate significantly from the majority of nodes. This paper targets node-level fraud detection on static graphs.
GraphConsis(Liu et al. 2020) combines contextual embeddings with node features and introduces node-related relationship attention weights. CARE-GNN(Dou et al. 2020) aims at the camouflage problem in fraud detection, and designs GNN enhanced neighbor selection module. PC-GNN(Liu et al. 2021b) designs adaptive subgraph and neighbor sampling methods to fraud detection scenarios. GAGA(Wang et al. 2023) uses group aggregation to generate discriminative neighborhood information to handle low homophily problems and enhances the original feature space with learnable encoding. This brings some inspiration to our work. GTAN(Xiang et al. 2023) designs gated temporal attention networks to learn node transaction representations.
@@
GPRGNN(Chien et al. 2021) combines Adaptive Generalized PageRank (GPR) Schemes with GNNs to adapt to heterophily graphs. ACM(Luan et al. 2022) studies heterophily from the perspective of node similarity after aggregation, and define a new homophily index. H2-FDetector(Shi et al. 2022) identifies homophily and heterophily connections through known labeled nodes, and designs different aggregation strategies for them. BWGNN(Tang et al. 2022) designs a band-pass filter kernel function to transmit the information of each frequency band separately, thus ensure that the spectrum energy distribution is consistent before and after message delivery. GHRN(Gao et al. 2023a) judges and cuts heterophily edges from the perspective of graph spectrum.
@@
We first introduce the notations of this paper, and then give the formal problem definition. Let $G=(\mathcal{V},\mathcal{E},X,Y)$ denotes an attributed fraud graph, where $\mathcal{V}=\{v_{1},\cdots,v_{N}\}$ denotes its node set, $\mathcal{E}$ is the edge set, and $A\in\{0,1\}^{N\times N}$ is the adjacency matrix of $G$ with $A_{ij}=1$ if there is a connection between node $v_{i}$ and $v_{j}$. $X\in R^{N\times d}$ denotes the matrix of node attributes where $x_{i}\in R^{d}$ is the attribute vector of $v_{i}$, and $Y$ is the corresponding labels of nodes.
We always formulate the graph-based fraud detection(GFD) as a semi-supervised node-level binary classification problem. Most of the time, anomalies are regarded as positive with label 1, while normal nodes are seen as negative with label 0(Jing et al. 2021). And nodes of the whole fraud graph are divided into the labeled ones and the unlabeled ones, the former are labeled with $Y_{train}$ and the labels of latter ones $Y_{test}$ are invisible during training period. Thus GFD aims to learn an anomaly labeling function for our unlabeled nodes with known information.
@@
As $A$ is the adjacency matrix of $G$, we let the Graph Laplacian matrix $L$ be defined as $D-A$ or as $I-D^{-1/2}AD^{-1/2}$(Kipf and Welling 2016), where $I$ is an identify matrix. $L$ is a positive semidefinite matrix and can be rewritten as $U\Lambda U^{T}$, where $\Lambda$ is a diagonal matrix consisted of eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\ ...\leq\lambda_{N}$ and $U$ is composed of corresponding $N$ eigenvectors $U=(u_{1},u_{2},...,u_{N})$(Bo et al. 2021).
According to theory of graph signal processing, we can treat the eigenvectors of normalized Laplacian matrix $U$ as the bases in Graph Fourier Transform. Suppose $X=[x_{1},x_{2},...,x_{N}]$ are signals on the graph, we treat $U^{T}X$ as the graph fourier transform(Hoang, Maehara, and Murata 2021) of signal $X$. Thus, the convolution between the signal $x$ and convolution kernel $g$ is:
@@
where $\odot$ is the element-wise product of vectors and $g_{w}$ is a diagonal matrix representing the convolutional kernel in the spectral domain. Most popular subsequent spectral GNN methods mainly focused on improving the design of $g_{w}$. And for signals $X$, the Rayleigh quotient represents signals’ standardized variance fraction(Kim et al. 2015):
@@
where $|\mathcal{N}(v)|$ is the number of neighbours of node $v$, and $|{\mathcal{E}}|$ is the total number of edges in the whole graph, $y_{src}$ and $y_{dst}$ are the labels of the source and target node of the edge $\varepsilon$ respectively.
@@
Fig. 2 illustrates our framework. We propose a Spectrum-Enhanced and Environment-Constrainted Graph Fraud Detector(SEC-GFD) to comprehensively integrate rich spectrum and label information into a fraud detector. In particular, SEC-GFD first decompose the spectrum of the graph signal. Then it performs complicated message passing technique based on these frequency bands respectively to learn comprehensive representations. Then we design a local environment constraint module to iteratively enhance the high-order associations between nodes’ true label and their surrounding neighbor environments.
@@
In the process of designing a hybrid band-pass filter, For purpose of ensuring the division of the spectrum into appropriate frequency bands, we use BWGNN as our backbone. According to this, the beta wavelet transform kernel function can be written as:
@@
It is observed that $\mathcal{W}_{p,q}$ is actually a C power polynomial, and $C$ represents how many hops of neighbors’ information are considered. Since the Laplacian matrix measures the difference of nodes, it can overcome the over-smoothing problem of vanilla GNN to a certain extent and can be extended to higher order neighbors. When unpacking $\mathcal{W}_{p,q}$, $W_{C,0}$ denotes the C power of the Laplacian matrix $L$. The remained polynomials are the couplings of high-frequency information of neighbors below order C and low-frequency information of first-order neighbors. Furthermore, pure high-frequency information of low-order neighbors is lacked in these polynomials. Thus we expand corresponding part and design a hybrid-pass filter. When adding the pure high-frequency information from order 1 to $(C-1)$, we use $\varepsilon I-D^{-1/2}AD^{-1/2}$ (or $\varepsilon I-\tilde{A}$)(Bo et al. 2021) as the basis, where $\varepsilon$ is a scaling hyperparameter bounded in $[0,1]$. The low-order neighbor high-frequency filtering transformation is expressed as follows:
$\begin{split}R_{1}:\mathcal{F}_{H}^{1}&=\varepsilon I-\tilde{A}=(\varepsilon-1)I+L,\\ R_{2}:\mathcal{F}_{H}^{2}&=(\varepsilon I-\tilde{A})^{2}=((\varepsilon-1)I+L)^{2},\\ \vdots\\ R_{C-1}:\mathcal{F}_{H}^{C-1}&=(\varepsilon I-\tilde{A})^{C-1}=((\varepsilon-1)I+L)^{C-1}\\ \end{split}$ (7)
@@
$\begin{split}H_{0}&=MLP(X),\\ \mathcal{B}_{i}&=W_{i,C-i}H_{0}\quad\mathcal{H}_{j}=R_{j}H_{0},\\ H&=f_{agg}[\mathcal{B}_{0},\cdotp\cdotp\cdotp,\mathcal{B}_{C},\mathcal{H}_{1},\cdotp\cdotp\cdotp,\mathcal{H}_{C-1}],\end{split}$ (9)
where $MLP$(Taud and Mas 2018) denotes a simple multi-layer perceptron and $f_{agg}$ can be any aggregation function such as summation or concatenation. $H$ is the aggregated representation of hybrid-pass filter, then weighted cross-entropy loss(Zhong, Wang, and Miao 2019) is used for the training of Hybrid-Pass Filter Module:
@@
So far, we have obtained the multi-hop neighbor information of the target nodes and their k-nearest neighbor environmental feature information. In order to realize the local environmental constraints, this paper judges based on common senses that, the multi-hop neighbor features learned by normal nodes and the local environmental features should be as similar as possible, while the two features learned by anomaly nodes should have a large deviation. Following the example of InfoNCE loss(You et al. 2020) in Contrastive Learning, we designed a constraint loss function $\mathcal{L}_{env}$ specially adapted to the above judgments:
@@
where $\mathcal{V}_{n}$ and $\mathcal{V}_{a}$ represent the sets of normal nodes and anomaly nodes in the training set respectively, and $sim(\cdot,\cdot)$ is the simple cosine similarity operator.
@@
$\mathcal{L}=\alpha\mathcal{L}_{hybrid}+(1-\alpha)\mathcal{L}_{env},$ (15)
@@
In this section, we evaluate the performance of our proposed model on real-world datasets through a series of experiments, and compare the results with those of state-of-the-art baselines to demonstrate the effectiveness of our approach. In addition, we conduct experiments on the response of heterophily edges to distinct filters, as well as experiments on ablation and hyperparameter analysis.
@@
Dataset. We conduct experiments on four real-world datasets targeted at fraud detection scenarios. Overall, they are Amazon(McAuley and Leskovec 2013), YelpChi(Rayana and Akoglu 2015), and two recently published transaction datasets, i.e., T-Finance and T-Social(Tang et al. 2022). The YelpChi dataset consists of filtered and recommended hotel and restaurant reviews from Yelp. The Amazon dataset includes product reviews in the music instrument category. The T-Finance dataset and T-Social dataset detect anomalous accounts in transaction networks and social networks respectively, they share the same 10-dimensional features. And the T-Social dataset is a large-scale dataset, several tens to hundreds of times larger in scale compared to the other datasets. The statistics of these datasets are summarized in Table 1.
Table 1: Statistics of Datasets
Dataset Nodes Edges
Anomaly(%)
Features
Amazon 11944 4,398,392
6.87
25
YelpChi 45954 3,846,979
14.53
32
T-Finance 39357 21,222,543
4.58
10
T-Social 5,781,065 73,105,508
3.01
10
Table 2: Fraud Detection Performance on Threat Datasets
@@
F1-macro
AUC
F1-macro
AUC
F1-macro
AUC
F1-macro
AUC
Homophily GNN GCN
53.21
56.80
63.52
80.14
71.47
66.31
61.73
87.84
GraphSAGE
64.34
73.61
75.25
86.73
55.27
69.52
60.49
72.65
GAT
55.86
58.92
82.41
89.73
54.27
75.29
73.72
88.61
GIN
63.42
75.28
71.14
81.38
65.72
81.63
63.38
81.33
GFD Algorithm GraphConsis
57.91
69.55
78.46
87.27
73.58
91.42
58.32
73.89
CARE-GNN
63.58
79.42
83.31
91.57
65.37
91.93
55.92
68.78
PC-GNN
64.75
77.14
91.73
95.63
77.25
92.83
57.21
71.27
GAGA 76.71
89.77
90.31
95.61
85.13
94.72
76.64
87.93
Heterophily GNN ACM
69.72
88.28
81.83
93.69
85.48
96.02
81.27
90.79
H2-FDetector
69.38
88.63
83.84
96.41
87.39
95.41
78.89
88.56
BWGNN
76.66
90.56
91.68 97.52
84.59
95.29
84.58 95.27
GDN
76.05
90.79
90.68
97.09
86.12
94.28
80.63
89.35

### v3

@@ Abstract
Graph-based fraud detection (GFD) can be regarded as a challenging semi-supervised node binary classification task. In recent years, Graph Neural Networks (GNN) have been widely applied to GFD, characterizing the anomalous possibility of a node by aggregating neighbor information. However, fraud graphs are inherently heterophilic, thus most of GNNs perform poorly due to their assumption of homophily. In addition, due to the existence of heterophily and class imbalance problem, the existing models do not fully utilize the precious node label information. To address the above issues, this paper proposes a semi-supervised GNN-based fraud detector SEC-GFD. This detector includes a hybrid filtering module and a local environmental constraint module, the two modules are utilized to solve heterophily and label utilization problem respectively. The first module starts from the perspective of the spectral domain, and solves the heterophily problem to a certain extent. Specifically, it divides the spectrum into various mixed-frequency bands based on the correlation between spectrum energy distribution and heterophily. Then in order to make full use of the node label information, a local environmental constraint module is adaptively designed. The comprehensive experimental results on four real-world fraud detection datasets denote that SEC-GFD outperforms other competitive graph-based fraud detectors. We release our code at https://github.com/Sunxkissed/SEC-GFD.
@@ Introduction
With the rapid development of the Internet, various fraudulent entities emerge in fraud activities, and the huge losses caused by this have attracted continuous attention from industry, academia and state agencies (Wang et al. 2018; Pourhabibi et al. 2020; Liu et al. 2021a; Zhang et al. 2021). For instance, fraudulent behaviors aimed at online payment have caused huge property losses to clients; and fraudulent or malicious comments on websites may bring negative effects on merchants and consumers.
@@
In recent years, graph-based fraud detection methods (Hooi et al. 2017; Qin et al. 2022) have been widely used in practical applications. Using entities in fraud scenarios as nodes and interactions between entities as edges, graph-level fraud detection can be achieved. Furthermore, based on the constructed fraud graph, as a deep neural network for graph representation learning, Graph Neural Network (GNN) (Jin et al. 2021) has been demonstrated to be effective in this task. Specifically, GNN employs neural modules to aggregate neighbor information and update node representations, enabling the discrimination of fraudulent entities.
@@
However, fraud detection scenarios are inherently heterophily (Zhu et al. 2021), i.e., target nodes (like node 5 in Figure 1) and their direct neighbors are prone to different class labels or features. In practical applications, fraudulent entities may use benign entities as springboards to perform high-order interactions with multi-hop neighbors to accomplish fraudulent actions or transmit fraudulent information, e.g., fraud node 3 and 6 are two-hop neighbors in Figure 1. For instance, in financial transaction networks (Akoglu, Tong, and Koutra 2015), fraudsters often use non-fraudulent users to conduct transactions.
@@
Due to the existence of heterophily, the performance of conventional graph neural network is unsatisfactory. This is because vanilla GNN is naturally a low-pass filter, which forces the representation of adjacent nodes to be similar during the information aggregation process. Then it may destroy the discriminative information of anomaly nodes. Furthermore, in the domain of fraud detection, heterophily graph datasets exhibit highly imbalanced class distribution properties. This phenomenon results in anomaly nodes being submerged in normal nodes, which manifests as high homophily of normal nodes and high heterophily of anomaly nodes. This is also called the common camouflage problem (Zhang et al. 2023) in fraud detection scenarios. Moreover, current semi-supervised fraud detection methods based on Graph Neural Networks (GNNs) suffer from limitations in terms of label utilization. These approaches typically learn node representations through a message passing mechanism, extract node embeddings from the final layer of the network, and incorporate label information during the training process. Evidently, these approaches overlook the intricate high-order associations between node label information, node characteristics, and their contextual surroundings.
@@
The current paramount concern entails addressing the heterophily issue within the context of extreme category imbalance. The advanced methods can be divided into two types. The first type of the methods is to design a new message aggregation mechanism (Xu et al. 2023) to simultaneously or separately consider the characteristics of homophily and heterophily edges. The second type of the methods is to identify the heterophily edges on fraud graphs and then perform pruning. According to recent researches, scholars found the spectral ‘right-shift’ phenomenon (Tang et al. 2022) on both synthetic and real datasets, proving that the higher the anomaly ratio, the more high-frequency components on the graph spectrum. In addition, researchers also demonstrate that, when using the ground truth labels as signals, the proportion of heterophily edges is positively correlated with the spectral energy distribution (Gao et al. 2023a). Based on these, we conclude that heterophily is not universally harmful, instead it ought to respond distinctively when meeting different filters. In order to verify our conjecture, we conduct sufficient experiments, and the specific related discussions are placed in experiments.
Figure 1: Schematic diagram of fraud relationship in real scenarios.
@@
In order to design a GNN-based fraud detector that adapts to high heterophily and has remarkable label utilization, we first propose a conjecture that heterophily edge clipping responds differently to different filters, and then conduct comprehensive experiments to verify the conjecture. Based on the results, we design a Spectrum-Enhanced and Environment-Constrainted Graph Fraud Detector (SEC-GFD) to comprehensively integrate rich spectrum and label information into a fraud detector. The framework of the proposed method is shown in Figure 2, consisting of two different modules that address the aforementioned two challenges respectively. The hybrid-pass filter module employs a sophisticated message passing technique utilizing the decomposed spectrum to acquire comprehensive high-frequency representations. The local environmental constraint module enhances the higher-order connections between target nodes and their surrounding environments, thereby mitigating the issue of limited label utilization. The main contributions of this method are summarized as follows:
@@
  • To address the challenge of heterophily, this paper employs spectrum analysis to partition the graph spectrum into hybrid frequency bands. Then a novel hybrid spectral filter is designed to conduct message passing on these frequency bands separately and finally aggregate.
@@
Semi-supervised node classification is one of the most important tasks in graph learning. In recent years, GNNs have achieved excellent performance in semi-supervised node classification, like GCN (Kipf and Welling 2016), GraphSAGE (Hamilton, Ying, and Leskovec 2017), GAT (Velivcković et al. 2017), and GIN (Xu et al. 2018). Compared with traditional graph embedding methods such as DeepWalk (Perozzi, Al-Rfou, and Skiena 2014) and Node2vec (Grover and Leskovec 2016), which mainly focus on graph structure, GNN can consider both node features and graph structure information. Each layer of GNN first performs message passing according to the node neighbors, and then updates the node representation through the aggregated information.
Figure 2: The above image presents an overview of our model SEC-GFD, where (a) and (b) respectively demonstrate the details of the hybrid-pass filter module and local environmental constraint module.
@@
GNN-based fraud detection (GFD) can be defined as an imbalanced binary classification task that identifies outlier nodes deviating significantly from the majority of nodes. This paper targets node-level fraud detection on static graphs. For instance, GraphConsis (Liu et al. 2020) combines contextual embeddings with node features and introduces node-related relationship attention weights. CARE-GNN (Dou et al. 2020) aims at the camouflage problem in fraud detection, and designs GNN enhanced neighbor selection module. PC-GNN (Liu et al. 2021b) designs adaptive subgraph and neighbor sampling methods to fraud detection scenarios. GAGA (Wang et al. 2023) uses group aggregation to generate discriminative neighborhood information to handle low homophily problems and enhances the original feature space with learnable encoding. This brings some inspiration to our work. GTAN (Xiang et al. 2023) designs gated temporal attention networks to learn node transaction representations.
@@
For instance, GPRGNN (Chien et al. 2020) combines adaptive Generalized PageRank (GPR) schemes with GNNs to adapt to heterophily graphs. ACM (Luan et al. 2022) studies heterophily from the perspective of node similarity after aggregation, and then defines a new homophily index. H2-FDetector (Shi et al. 2022) identifies homophily and heterophily connections through known labeled nodes, and designs different aggregation strategies for them. BWGNN (Tang et al. 2022) designs a band-pass filter kernel function to transmit the information of each frequency band separately, thus ensuring that the spectrum energy distribution is consistent before and after message passing. GHRN (Gao et al. 2023a) judges and cuts heterophily edges from the perspective of graph spectrum.
@@
We first introduce the notations of this paper, and then give the formal problem definition. Let $G=(\mathcal{V},\mathcal{E},X,Y)$ denote an attributed fraud graph, where $\mathcal{V}=\{v_{1},\cdots,v_{N}\}$ denotes its node set, $\mathcal{E}$ is the edge set, and $A\in\{0,1\}^{N\times N}$ is the adjacency matrix of $G$ with $A_{ij}=1$ if there is a connection between node $v_{i}$ and $v_{j}$. $X\in R^{N\times d}$ denotes the matrix of node attributes where $x_{i}\in R^{d}$ is the attribute vector of $v_{i}$, and $Y$ is the corresponding labels of the nodes.
We always formulate the graph-based fraud detection (GFD) as a semi-supervised node-level binary classification problem. Most of the time, anomalies are regarded as positive with label 1, while normal nodes are viewed as negative with label 0 (Jing et al. 2021). The nodes of the whole fraud graph are divided into the labeled ones and the unlabeled ones, the former are labeled with $Y_{train}$ and the labels of latter ones $Y_{test}$ are invisible during training period. Thus GFD aims to learn an anomaly labeling function for our unlabeled nodes with all known information:
@@
As $A$ is the adjacency matrix of $G$, we let the graph laplacian matrix $L$ be defined as $D-A$ or as $I-D^{-1/2}AD^{-1/2}$ (Kipf and Welling 2016), where $I$ is an identify matrix. $L$ is a positive semidefinite matrix and can be rewritten as $U\Lambda U^{T}$, where $\Lambda$ is a diagonal matrix consisted of eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\ ...\leq\lambda_{N}$ and $U$ is composed of the corresponding $N$ eigenvectors $U=(u_{1},u_{2},...,u_{N})$ (Bo et al. 2021).
According to theory of graph signal processing, we can treat the eigenvectors of normalized Laplacian matrix $U$ as the bases in Graph Fourier Transform. Suppose $X=[x_{1},x_{2},...,x_{N}]$ are signals on the graph, we treat $U^{T}X$ as the graph fourier transform (Hoang, Maehara, and Murata 2021) of signal $X$. Thus, the convolution between the signal $x$ and convolution kernel $g$ is as follows:
@@
where $\odot$ is the element-wise product of vectors and $g_{w}$ is a diagonal matrix representing the convolutional kernel in the spectral domain. Most popular subsequent spectral GNN methods mainly focus on improving the design of $g_{w}$. And for signals $X$, the Rayleigh quotient represents signals’ standardized variance fraction (Kim et al. 2015):
@@
where $|\mathcal{N}(v)|$ is the number of neighbours of node $v$, and $|{\mathcal{E}}|$ is the total number of edges in the whole graph, $y_{src}$ and $y_{dst}$ are the labels of the source and destination node of the edge $\varepsilon$ respectively.
@@
Figure 2 illustrates the details of our framework. We propose a Spectrum-Enhanced and Environment-Constrainted Graph Fraud Detector (SEC-GFD) to comprehensively integrate rich spectrum and label information into a fraud detector. In particular, SEC-GFD first decomposes the spectrum of the graph signal. Then it performs complicated message passing technique based on these frequency bands respectively to learn comprehensive representations. Then we design a local environmental constraint module to iteratively enhance the high-order associations between nodes’ true labels and their surrounding neighbor environments.
@@
In the process of designing a hybrid band-pass filter, for purpose of ensuring the division of the spectrum into appropriate frequency bands, we utilize BWGNN (Tang et al. 2022) as our backbone. According to this, the beta wavelet transform kernel function can be written as:
@@
It is observed that $\mathcal{W}_{p,q}$ is actually a C power polynomial, and $C$ represents how many hops of neighbors’ information are considered. Since the Laplacian matrix measures the differences of nodes, it can overcome the over-smoothing problem of vanilla GNN to a certain extent and can be extended to higher order neighbors. When unpacking $\mathcal{W}_{p,q}$, $W_{C,0}$ denotes the C power of the Laplacian matrix $L$. The remained polynomials are the couplings of high-frequency information of neighbors below order C and low-frequency information of low-order neighbors. Furthermore, pure high-frequency information of low-order neighbors is lacked in these polynomials, thus we expand the corresponding part and design a hybrid-pass filter. When adding the pure high-frequency information from order 1 to $(C-1)$, we use $\varepsilon I-D^{-1/2}AD^{-1/2}$ or $\varepsilon I-\tilde{A}$ (Bo et al. 2021) as the basis, where $\varepsilon$ is a scaling hyperparameter bounded in $[0,1]$. The low-order neighbor high-frequency filtering transformation is expressed as follows:
$\begin{split}R_{1}:\mathcal{F}_{H}^{1}&=\varepsilon I-\tilde{A}=(\varepsilon-1)I+L,\\ R_{2}:\mathcal{F}_{H}^{2}&=(\varepsilon I-\tilde{A})^{2}=((\varepsilon-1)I+L)^{2},\\ \vdots\\ R_{C-1}:\mathcal{F}_{H}^{C-1}&=(\varepsilon I-\tilde{A})^{C-1}=((\varepsilon-1)I+L)^{C-1}.\\ \end{split}$ (7)
@@
$\begin{split}H_{0}&=MLP(X),\\ \mathcal{B}_{i}&=W_{i,C-i}H_{0}\quad\mathcal{H}_{j}=R_{j}H_{0},\\ H&=f_{agg}(\mathcal{B}_{0},\cdotp\cdotp\cdotp,\mathcal{B}_{C},\mathcal{H}_{1},\cdotp\cdotp\cdotp,\mathcal{H}_{C-1}),\end{split}$ (9)
where $MLP$ (Taud and Mas 2018) denotes a simple multi-layer perceptron and $f_{agg}$ can be any aggregation function (Wu et al. 2023) such as summation or concatenation. $H$ is the aggregated representation of hybrid-pass filter, then weighted cross-entropy loss (Zhong, Wang, and Miao 2019) is used for the training of hybrid-pass filter module:
@@
So far, we have obtained the multi-hop neighbor information of the target nodes and their k-nearest neighbor environmental feature information. In order to realize the local environmental constraints, this paper judges based on common senses that, the multi-hop neighbor features learned by normal nodes and the local environmental features should be as similar as possible, while the two features learned by anomaly nodes should have a large deviation. Following the formula of InfoNCE loss (You et al. 2020) in Contrastive Learning, we design a constraint loss function $\mathcal{L}_{env}$ specially adapted to the above judgments:
@@
where $\mathcal{V}_{n}$ and $\mathcal{V}_{a}$ represent the set of normal nodes and anomaly nodes in the training set respectively, and $sim(\cdot,\cdot)$ is the simple cosine similarity operator.
@@
$\mathcal{L}=\alpha\mathcal{L}_{hybrid}+(1-\alpha)\mathcal{L}_{env}.$ (15)
@@
In this section, we evaluate the performance of our proposed approach on real-world datasets through a series of experiments, and compare the results with those of state-of-the-art baselines to demonstrate the effectiveness of our approach. In addition, we conduct experiments on the response of heterophily edges to distinct filters, as well as experiments on ablation and hyperparameter analysis.
@@
Dataset. We conduct experiments on four real-world datasets targeted at fraud detection scenarios. Overall, they are Amazon (McAuley and Leskovec 2013), YelpChi (Rayana and Akoglu 2015), and two recently published transaction datasets, i.e., T-Finance and T-Social (Tang et al. 2022). The YelpChi dataset consists of filtered and recommended hotel and restaurant reviews from Yelp. The Amazon dataset includes product reviews in the music instrument category. The T-Finance dataset and T-Social dataset detect anomalous accounts in transaction networks and social networks respectively, and they share the same 10-dimensional features. The T-Social dataset is a large-scale dataset, several tens to hundreds of times larger in scale compared to the other datasets. The statistics of these datasets are summarized in Table 1.
Dataset Nodes Edges Anomaly(%) Features
Amazon 11,944 4,398,392 6.87 25
YelpChi 45,954 3,846,979 14.53 32
T-Finance 39,357 21,222,543 4.58 10
T-Social 5,781,065 73,105,508 3.01 10
Table 1: Statistics of four datasets.
@@
F1-macro AUC F1-macro AUC F1-macro AUC F1-macro AUC
Homophily GNN GCN 53.21 56.80 63.52 80.14 71.47 66.31 61.73 87.84
GraphSAGE 64.34 73.61 75.25 86.73 55.27 69.52 60.49 72.65
GAT 55.86 58.92 82.41 89.73 54.27 75.29 73.72 88.61
GIN 63.42 75.28 71.14 81.38 65.72 81.63 63.38 81.33
GFD Algorithm GraphConsis 57.91 69.55 78.46 87.27 73.58 91.42 58.32 73.89
CARE-GNN 63.58 79.42 83.31 91.57 65.37 91.93 55.92 68.78
PC-GNN 64.75 77.14 91.73 95.63 77.25 92.83 57.21 71.27
GAGA 76.71 89.77 90.31 95.61 85.13 94.72 76.64 87.93
Heterophily GNN ACM 69.72 88.28 81.83 93.69 85.48 96.02 81.27 90.79
H2-FDetector 69.38 88.63 83.84 96.41 87.39 95.41 78.89 88.56
BWGNN 76.66 90.56 91.68 97.52 84.59 95.29 84.58 95.27
GDN 76.05 90.79 90.68 97.09 86.12 94.28 80.63 89.35